Divide 34 pounds 8 ounces by 3
Answer by lenny460(1073)

Divide 34 pounds 8 ounces by 3.
34/3 = 11 pounds remainder 1 pound
1 pound = 16 ounces
16 + 8 = 24 ounces
24/3 = 8 ounces
The Answer:
11 pounds 8 ounces


obuong3@aol.com